# Photon transport and blockade based on non-Markovian interactions between a microring resonator and waveguide

### v1

@@ Abstract
We investigate photon transport and blockade based on the architecture where a waveguide is coupled to a microring resonator at two distinct points. This two-point coupling configuration between the waveguide and resonator gives rise to non-Markovian dynamics, which is induced by the photon transmission delay in the waveguide between the two coupled points. On one hand, by designing the non-Markovian coupling parameters between the waveguide and resonator, single- or two-photon transport and the resulting photon blockade effect can be manipulated according to the output photonic states at the end of waveguide. This non-Markovian process can be evaluated by scattering matrices and second order correlation functions related to the distance between two coupled points. On the other hand, when classical driving fields are applied upon the resonator with interactions between its clockwise and counterclockwise modes, the blockade effect of the output field can be determined by the intracavtiy eigenstates. Then the correlations of the output field as well as the intracavity states can be modulated by the non-Markovian coupling between the microring resonator and waveguide.
@@ I Introduction
The generation of single-photon states is a crucial aspect in the realization of quantum information processing based on photonic integrated circuits [16]. Single-photon states can be generated through spontaneous emission when a quantum emitter is in free space [17], coupled to a waveguide [18], within a cavity [19], or modulated by nonlinear materials [20]. The photon blockade effect, which can suppress multiphoton events by blocking the generation of a second photon, has been widely applied to enhance the reliability of on-demand single-photon sources [21, 22, 23, 24]. For instance, in the experimental realization of low-temperature atom-cavity systems, the significant nonlinear interactions between the atom and a Fabry-Perot cavity can induce the blockade effect, thereby influencing the transmission of photons in the cavity [25]. Consequently, there can be more or less than one photon detected at the output side of the cavity. This experimental setup can be similarly generalized to the platforms based on microing resonators, which are commonly employed in the room-temperature quantum information processing [26, 27]. Similar to a Fabry-Perot cavity, a microring resonator can be coupled to atoms [28, 29, 30, 31], waveguides used for the mediation of photons [32, 33], and Kerr materials with nonlinearity [34, 35, 36], resulting in linear and nonlinear quantum dynamics [37, 38, 39].
@@
Moreover, non-Markovian quantum photonics can arise in systems where there are multiple microring resonators coupled to a waveguide or single microring resonator coupled to a waveguide at different points, due to the transmission delays of photons in the waveguide [40]. For example, a microring resonator can be coupled to a waveguide via multi-point coupling, which contrasts with the traditional single-point coupling configuration [39]. For another case, one resonator can be coupled to a waveguide at two distinct points, then the photons transmitted between two coupled points can construct a coherent feedback channel [41]. In these non-Markovian networks based on resonator-waveguide interactions, the photon transmission delay in the waveguide can influence the resonator dynamics [16], as well as the photon transport properties in the waveguide [42, 34, 43].
@@
In this paper, we investigate photon transport and blockade based on the non-Markovian interaction realized by the configuration where a microring resonator is coupled to a waveguide at two points. The photon transport and blockade can be manipulated by the coupling strength between waveguide and resonator, the transmission delay of photons in the waveguide between the two coupled points, and the nonlinear coupling between the resonator and Kerr material. The remainder of this paper is organized as follows. In Sec. II, we derive the non-Markovian interaction model between the microring resonator and waveguide as a non-Markovian input-output formalism. Based on this, in Sec. III, we study the non-Markovian transport of single and two photons evaluated by the scattering matrix, as well as the possibility of the occurrence of photon blockade. In Sec. IV, we study the non-Markovian dynamics when the resonator is driven by classical control fields, and the resulting blockade effect related with the two-, single- and zero-photon states. Finally, conclusions are presented in Sec. V.
@@ II Non-Markovian interactions between microring resonator and waveguide
$+\int_{-\infty}^{\infty}\left[ie^{\frac{i\omega x}{c}}\left(\sqrt{\frac{g_{1}}{2\pi}}+\sqrt{\frac{g_{2}}{2\pi}}e^{-\frac{i\omega L}{c}}\right)b^{{\dagger}}_{\omega}a+{\rm H.c.}\right]\mathrm{d}\omega$
$+\int_{-\infty}^{\infty}\left[ie^{-\frac{i\omega x}{c}}\left(\sqrt{\frac{\tilde{g}_{1}}{2\pi}}+\sqrt{\frac{\tilde{g}_{2}}{2\pi}}e^{\frac{i\omega L}{c}}\right)c^{{\dagger}}_{\omega}\tilde{a}+{\rm H.c.}\right]\mathrm{d}\omega,$
@@
Apart from the non-Markovian interactions between the microring resonator and waveguide, the photon transport in Fig. 1 is also influenced by the resonator’s loss rate, and the coupling strengths among the Kerr material, clockwise and counterclockwise resonator modes. Based on these interactions concluded in Eq. (5), we can derive the non-Markovian input-output formalism as follows [49, 42].
@@ III.2 Two-photon transport and blockade
When there are two input photons, the two-photon transport can be affected by the Kerr nonlinearity due to the Hamiltonian in Eq. (3), which is different from the single-photon circumstance [34]. Generalized from the above single-photon case, in the two-photon transport, we denote the input photon frequencies as $\omega_{1}$ and $\omega_{2}$, and the output photon frequencies as $\nu_{1}$ and $\nu_{2}$. Thus be taking $N=2$ in Eq. (15), the two-photon scattering matrix reads [42]
@@
where $t_{\nu_{j}}$ with $j=1,2$ represents the single-photon process by taking $\omega=\nu_{j}$ in Eq. (25),
@@
with $\omega=\nu_{j}$ has a similar format as $t_{\omega}$, we additionally define the following function for the input and output photon frequencies
$\Gamma(\omega)=\frac{\sqrt{g_{1}}+\sqrt{g_{2}}e^{i\omega\frac{L}{c}}}{\kappa_{a}+i\left(\omega_{a}-\omega\right)},$ (29)
and the detailed derivations are given in Appendix B.
@@
When there are two input photons with frequencies $\omega_{1}$ and $\omega_{2}$, we denote the input photon state as $\left|\psi_{\rm in}\right\rangle$ with the amplitude $\alpha$ in the format of a coherent state as [51]
$\left|\psi_{\rm in}\right\rangle=\left|\alpha_{1}\right\rangle\left|\alpha_{2}\right\rangle,$ (32)
and we assume that the two photons have the same amplitude $\alpha$, then combined with the format of $b_{\rm out}(t)$ in Eq. (23), one of the denominator in Eq. (31) for an input photon mode $\omega$ reads [51]
@@
Figure 3: Second order correlation function $\ln\left[g_{\rm out}^{(2)}(0)\right]$ of the output field is influenced by input photon frequencies and other parameters when $\kappa_{a}=1$MHz and $\chi=0.1$MHz. In (a) and (b), $g_{1}=g_{2}=1$MHz. In (c) and (d), $g_{2}=2g_{1}=4/3$MHz. In (a) and (c), $\omega_{a}L/c=\pi$. In (b) and (d), $\omega_{1}+\omega_{2}=2\omega_{a}$.
As compared in Fig. 3, the occurrence of photon blockade is influenced by the coupling strength between the resonator and waveguide, the frequencies of two input photons, and the non-Markovian interaction evaluated by the length of the waveguide $L$ between two coupled points. For example, the comparison between Figs. 3(a,c) for two uncorrelated input photons shows that the values of $g_{1}$ and $g_{2}$ can influence the occurrence of photon blockade. Due to the format of scattering matrix in the second line of Eq. (27), we take $\omega_{1}+\omega_{2}=2\omega_{a}$ in the simulations of Figs. 3(b,d). For the correlated input photons, the second order correlation functions of the output field can be influenced by both $L$ and the frequency of each input photon. The simulations in Fig. 3(d) also show that, the diverse between $g_{1}$ and $g_{2}$ can induce blockade effect when the frequencies of two input correlated photons are different.
@@ IV Non-Markovian photon blockade with a driven resonator
$+\left[\frac{g_{1}+g_{2}}{2}+\sqrt{g_{1}g_{2}}\cos\left(\frac{\omega_{a}L}{c}\right)+2\kappa_{a}\right]\mathcal{L}_{a}[\rho]$
$+\left[\frac{\tilde{g}_{1}+\tilde{g}_{2}}{2}+\sqrt{\tilde{g}_{1}\tilde{g}_{2}}\cos\left(\frac{\omega_{a}L}{c}\right)+2\kappa_{a}\right]\mathcal{L}_{\tilde{a}}[\rho],$
where $H_{a,\tilde{a}}$ represents the free Hamiltonian of the microring resonator independent from its interaction with the waveguide, and $\mathcal{L}_{O}[\rho]=O\rho O^{{\dagger}}-\left(\rho O^{{\dagger}}O+O^{{\dagger}}O\rho\right)/2$ for an arbitrary operator $O$. The master equation (37) can be similarly derived via the approach in Refs. [53, 59] or by replacing the atomic operator in Refs. [57, 58] with resonator operators, and more details are omitted.
@@
$g_{\rm eff}^{-}=\frac{g_{1}+g_{2}}{2}+\sqrt{g_{1}g_{2}}\cos\left(\frac{\omega_{a}L}{c}\right)+2\kappa_{a}\geq 0,$ (38a)
$g_{\rm eff}^{+}=\frac{\tilde{g}_{1}+\tilde{g}_{2}}{2}+\sqrt{\tilde{g}_{1}\tilde{g}_{2}}\cos\left(\frac{\omega_{a}L}{c}\right)+2\kappa_{a}\geq 0,$ (38b)
@@
$+\left(\hat{g}_{\rm eff}^{-}-i\frac{g_{\rm eff}^{-}}{2}\right)a^{{\dagger}}a+\left(\hat{g}_{\rm eff}^{+}-i\frac{g_{\rm eff}^{+}}{2}\right)\tilde{a}^{{\dagger}}\tilde{a}$.
@@
$~~~~~~~~-\left(\frac{g_{\rm eff}^{-}}{2}+i\hat{g}_{\rm eff}^{-}\right)C_{10},$ (43b)
$\dot{C}_{01}=-i\epsilon C_{11}-i\zeta^{*}C_{10}-\left(\frac{g_{\rm eff}^{+}}{2}+i\hat{g}_{\rm eff}^{+}\right)C_{01},$ (43c)
@@
$~~~~~~~~-\left(\frac{g_{\rm eff}^{-}}{2}+\frac{g_{\rm eff}^{+}}{2}+i\hat{g}_{\rm eff}^{-}+i\hat{g}_{\rm eff}^{+}+2i\chi\right)C_{11},$ (43d)
$\dot{C}_{02}=-i\sqrt{2}\zeta^{*}C_{11}-\left(g_{\rm eff}^{+}+2i\hat{g}_{\rm eff}^{+}+2i\chi\right)C_{02},$ (43e)
@@
Figure 4: Second order correlation function $\ln\left[g_{\rm out}^{(2)}(0)\right]$ influenced by Kerr nonlinearity when $g_{0}=1$MHz, $\kappa_{a}=0.2$MHz and $\zeta=0.5$MHz.
@@
Figure 5: The photon states in the resonator with $g_{0}=1$MHz, $\chi=4$MHz, $\kappa_{a}=0.2$MHz and $\zeta=0.5$MHz. As for the non-Markoivan parameter settings, $\omega_{a}L/c=\pi/3$ in (a) and (c), while $\omega_{a}L/c=\pi$ in (b) and (d). As for the classical driving field, $\epsilon=1$MHz in (a) and (b), while $\epsilon=4$MHz in (c) and (d).
To further clarify the photon number states in the microring resonator, the amplitudes of quantum states are simulated in Fig. 5. When the drive field is weak as $\epsilon=1$MHz, as in Fig. 5(a) with $\omega_{a}L=\pi/3$, the effective coupling between the waveguide and resonator is nonzero, which means that the photons in the resonator can be transmitted into the waveguide. However, because the driving field is weak and $\kappa_{a}>0$, there are almost no two-photon states in the resonator. When $\omega_{a}L=\pi$ and $\epsilon=1$MHz in Fig. 5(b), the effective coupling between the waveguide and resonator is zero according to Eq. (38), then the drive applied upon the resonator can compensate its decaying to the environment, then there can be two-photon states in the resonator. Based on this, when the drive field is stronger, as in Figs. 5(c,d) with $\epsilon=4$MHz, the populations for two-photon states can be larger, and photon blockade can occur due to the area in Fig. 4(b) with $\ln\left[g_{\rm out}^{(2)}(0)\right]<0$. For example, according to the steady numerical results in Fig. 5(c), $g_{\rm out}^{(2)}(0)\approx 0.98<1$, while in Fig. 5(d), $g_{\rm out}^{(2)}(0)\approx 0.5<1$. Besides, the comparison between Figs. 5(c,d) also illustrates that the non-Markovian interaction between resonator and waveguide evaluated by $\omega_{a}L$ can influence the steady photonic states in the microring resonator. When $\omega_{a}L=\pi$ in Fig. 5(d), the counterclockwise photon mode cannot be easily decayed to the waveguide, then the population for the steady two-photon states can be larger than that in Fig. 5(c).

### v2

@@ Abstract
We investigate photon transport and blockade based on the architecture where a waveguide is coupled to a microring resonator at two distinct points. This two-point coupling configuration between the waveguide and resonator gives rise to non-Markovian dynamics, which is induced by the photon transmission delay in the waveguide between the two coupled points. On one hand, by designing the non-Markovian coupling parameters between the waveguide and resonator, single- or two-photon transport and the resulting photon blockade effect can be manipulated according to the output photonic states at the end of waveguide. We for the fist time evaluate the occurrence of photon blockade with scattering matrices and second order correlation functions based on this non-Markovian proposal related to the length of waveguide between two coupled points. On the other hand, when classical driving fields are applied upon the resonator with interactions between its clockwise and counterclockwise modes, the blockade effect of the output field can be determined by the intracavity eigenstates. Then the correlations of the output field as well as the intracavity states can be modulated by the non-Markovian coupling between the microring resonator and waveguide.
@@ I Introduction
The generation of single-photon states is a crucial aspect in the realization of quantum information processing based on photonic integrated circuits [20]. Single-photon states can be generated through spontaneous emission when a quantum emitter is in free space [31], coupled to a waveguide [80], within a cavity [46], or modulated by nonlinear materials [50]. The photon blockade effect, which can suppress multiphoton events by blocking the generation of a second photon, has been widely applied to enhance the reliability of on-demand single-photon sources [41, 69, 68, 57]. For instance, in the experimental realization of low-temperature atom-cavity systems, the significant nonlinear interactions between the atom and a Fabry-Perot cavity can induce the blockade effect, thereby influencing the transmission of photons in the cavity [5]. Consequently, there can be more or less than one photon detected at the output side of the cavity. This experimental setup can be similarly generalized to the platforms based on microring resonators, which are commonly employed in the room-temperature quantum information processing [29, 65]. Similar to a Fabry-Perot cavity, a microring resonator can be coupled to atoms [2, 15, 39, 75], waveguides used for the mediation of photons [59, 23], and Kerr materials with nonlinearity [7, 13, 76], resulting in linear and nonlinear quantum dynamics [44, 55, 36].
@@
Moreover, non-Markovian quantum photonics can arise in systems where there are multiple microring resonators coupled to a waveguide or single microring resonator coupled to a waveguide at different points, due to the transmission delays of photons in the waveguide [3]. For example, when one resonator is coupled to a waveguide at two distinct points, the photons transmitted between two coupled points can construct a coherent feedback channel [17, 83, 72]. In such non-Markovian networks, the two-point coupling between the waveguide and resonator is influenced by different photonic modes in the waveguide. For each single mode, the phase-modulated interactions between the resonator and waveguide is similar to the Markovian interactions between a resonator and environment [20]. In Markovian quantum systems, especially for the circumstance with more than one photon, a most precise approach to analyze the photon transportation and blockade is calculating the scattering matrices [19, 7, 74]. However, this approach has not be generalized to non-Markovian quantum networks with more than one coupling points between the resonator and waveguide.
@@
In this paper, we for the first time introduce the approach in Refs. [19, 7, 74] to a quantum system with non-Markovian properties, realized by a waveguide coupled to a microring resonator at two distinct points. The photon transport and blockade can be manipulated by coupling strengths between the waveguide and resonator, the transmission delay of photons in the waveguide between the two coupled points, and the nonlinear coupling between the resonator and Kerr materials. The remainder of this paper is organized as follows. In Sec. II, we derive the non-Markovian interaction model between the microring resonator and waveguide as a non-Markovian input-output formalism. Based on this, in Sec. III, we study the non-Markovian transport of single and two photons evaluated by the scattering matrix, as well as the possibility of the occurrence of photon blockade. In Sec. IV, we study the non-Markovian dynamics when the resonator is driven by classical control fields, and the resulting blockade effect related with the two-, single- and zero-photon states. Finally, conclusions are presented in Sec. V.
@@ II Non-Markovian interactions between microring resonator and waveguide
$+\int_{0}^{\infty}\left[ie^{\frac{i\omega x}{c}}\left(\sqrt{\frac{g_{1}}{2\pi}}+\sqrt{\frac{g_{2}}{2\pi}}e^{-\frac{i\omega L}{c}}\right)b^{{\dagger}}_{\omega}a+{\rm H.c.}\right]\mathrm{d}\omega$
$+\int_{0}^{\infty}\left[ie^{-\frac{i\omega x}{c}}\left(\sqrt{\frac{\tilde{g}_{1}}{2\pi}}+\sqrt{\frac{\tilde{g}_{2}}{2\pi}}e^{\frac{i\omega L}{c}}\right)c^{{\dagger}}_{\omega}\tilde{a}+{\rm H.c.}\right]\mathrm{d}\omega,$
@@
Apart from non-Markovian interactions between the microring resonator and waveguide, the photon transport in Fig. 1 is also influenced by the resonator’s loss rate, and the coupling strengths among the Kerr material, clockwise and counterclockwise resonator modes. Based on these interactions concluded in Eq. (5), we can derive the non-Markovian input-output formalism as follows [22, 19].
@@ III.1 Single-photon transport
Both Eq. (25) and the numerical simulations in Fig. 2 show that, the transport of a single photon can be modulated by the phase difference during its transmission between the two coupled points in the waveguide, namely the parameter $\omega L/c$ in Eq. (25), and this is similar to the classical circumstance that electric fields transports in a waveguide coupled to a mircoring resonator at two points [60].
@@ III.2 Two-photon transport and blockade
When there are two input photons, the two-photon transport can be affected by the Kerr nonlinearity due to the Hamiltonian in Eq. (3), which is different from the single-photon circumstance [7]. Generalized from the above single-photon case, in the two-photon transport, we denote the input photon frequencies as $\omega_{1}$ and $\omega_{2}$, and the output photon frequencies as $\nu_{1}$ and $\nu_{2}$. Thus by taking $N=2$ in Eq. (15), the two-photon scattering matrix reads [19]
@@
where $t_{\nu_{j}}$ with $j=1,2$ represents the single-photon process by taking $\omega=\nu_{j}$ in Eq. (25), namely
@@
which has a similar format as $t_{\omega}$. We additionally define the following function for the input and output photon frequencies
$\Gamma(\omega)=\frac{\sqrt{g_{1}}+\sqrt{g_{2}}e^{i\omega\frac{L}{c}}}{\kappa_{a}+i\left(\omega_{a}-\omega\right)}.$ (29)
More details on $\eta_{\omega}$ and $\Gamma(\omega)$ are given in Appendix B. The scattering matrix in Eq. (27) shows that, in this non-Markovian network with Kerr nonlinearity, the two-photon transport is not simply a combination of two single-photon transport processes, namely the components in the first line of the right hand side of Eq. (27). A more complex term is the correlated term in the second line of Eq. (27), and this is induced by the nonlinear interactions between the microring resonator and the Kerr material, which makes it different from the transport of classical fields.
@@
When there are two input photons with frequencies $\omega_{1}$ and $\omega_{2}$, we denote the input photon state as $\left|\psi_{\rm in}\right\rangle$ as a product state of two coherent states with the identical amplitude $\alpha$ in the format of a coherent state as [71]
$\left|\psi_{\rm in}\right\rangle=\left|\alpha_{1}\right\rangle\otimes\left|\alpha_{2}\right\rangle,$ (32)
where $\left|\alpha_{j}\right\rangle=|0\rangle+\alpha|1\rangle_{j}+\left(\alpha^{2}/\sqrt{2}\right)|2\rangle_{j}+O\left(\alpha^{3}\right)$ with the index $j=1,2$ representing the two input photons. Then combined with the format of $b_{\rm out}(t)$ in Eq. (23), one of the denominator in Eq. (31) for an input photon mode $\omega$ reads [71]
@@
Figure 3: Second order correlation function $\ln\left[g_{\rm out}^{(2)}(0)\right]$ of the output field is influenced by input photon frequencies and other parameters when $\kappa_{a}=1$GHz and $\chi=0.1$GHz. In (a) and (b), $g_{1}=g_{2}=1$GHz. In (c) and (d), $g_{2}=2g_{1}=4/3$GHz. In (a) and (c), $\omega_{a}L/c=\pi$. In (b) and (d), $\omega_{1}+\omega_{2}=2\omega_{a}$.
As compared in Fig. 3, the occurrence of photon blockade is influenced by the coupling strength between the resonator and waveguide, the frequencies of two input photons, and the non-Markovian interaction evaluated by the length of the waveguide $L$ between two coupled points. For example, the comparison between Figs. 3(a,c) for two input photons shows that the values of $g_{1}$ and $g_{2}$ can influence the occurrence of photon blockade. Due to the format of scattering matrix in the second line of Eq. (27), we take $\omega_{1}+\omega_{2}=2\omega_{a}$ in the simulations of Figs. 3(b,d). In this setting, the second order correlation functions of the output field can be influenced by both $L$ and the frequency of each input photon. The simulations in Fig. 3(d) also show that, the diverse between $g_{1}$ and $g_{2}$ can induce blockade effect when the frequencies of two input photons are different.
The above comparisons in Fig. 3 indicate that, photon blockade is influenced by the two-point coupling strengths between resonator and waveguide, the input photon frequencies and the length of waveguide between two coupled points. As determined by the scattering matrix in Eq. (27), because of the existence of Kerr nonlinearity $\chi$ and decaying rate to the environment $\kappa_{a}$, the photon blockade will not only occur when $\omega_{a}L/c=\mathbf{n}\pi$ with $\mathbf{n}=1,2,\cdots$, but also occur when $\omega_{a}L/c\neq\mathbf{n}\pi$, and the input photon frequencies do not equal $\omega_{a}$. This is supported by the numerical simulations in Figs. 3(a,c) that $\ln\left[g_{\rm out}^{(2)}(0)\right]<0$ when $\omega_{1}=\omega_{2}\neq\omega_{a}$, and especially in Fig. 3(d) that $\ln\left[g_{\rm out}^{(2)}(0)\right]<0$ when $\omega_{a}L/c\neq\mathbf{n}\pi$ and $\omega_{1},\omega_{2}\neq\omega_{a}$.
@@ IV Non-Markovian photon blockade with a driven resonator
$+\left[\frac{g_{1}+g_{2}}{2}+\sqrt{g_{1}g_{2}}\cos\left(\frac{\omega_{a}L}{c}\right)+\kappa_{a}\right]\mathcal{L}_{a}[\rho]$
$+\left[\frac{\tilde{g}_{1}+\tilde{g}_{2}}{2}+\sqrt{\tilde{g}_{1}\tilde{g}_{2}}\cos\left(\frac{\omega_{a}L}{c}\right)+\kappa_{a}\right]\mathcal{L}_{\tilde{a}}[\rho],$
where $H_{a,\tilde{a}}$ represents the free Hamiltonian of the microring resonator independent from its interaction with the waveguide, and $\mathcal{L}_{O}[\rho]=2O\rho O^{{\dagger}}-\rho O^{{\dagger}}O-O^{{\dagger}}O\rho$ for an arbitrary operator $O$. The master equation (37) can be similarly derived via the approach in Refs. [51, 43] or by replacing the atomic operator in Refs. [33, 64] with resonator operators, and more details are omitted.
@@
$g_{\rm eff}^{-}=\frac{g_{1}+g_{2}}{2}+\sqrt{g_{1}g_{2}}\cos\left(\frac{\omega_{a}L}{c}\right)+\kappa_{a}\geq 0,$ (38a)
$g_{\rm eff}^{+}=\frac{\tilde{g}_{1}+\tilde{g}_{2}}{2}+\sqrt{\tilde{g}_{1}\tilde{g}_{2}}\cos\left(\frac{\omega_{a}L}{c}\right)+\kappa_{a}\geq 0,$ (38b)
@@
$+\left(\hat{g}_{\rm eff}^{-}-ig_{\rm eff}^{-}\right)a^{{\dagger}}a+\left(\hat{g}_{\rm eff}^{+}-ig_{\rm eff}^{+}\right)\tilde{a}^{{\dagger}}\tilde{a}$.
@@
$~~~~~~~~-\left(g_{\rm eff}^{-}+i\hat{g}_{\rm eff}^{-}\right)C_{10},$ (43b)
$\dot{C}_{01}=-i\epsilon C_{11}-i\zeta^{*}C_{10}-\left(g_{\rm eff}^{+}+i\hat{g}_{\rm eff}^{+}\right)C_{01},$ (43c)
@@
$~~~~~~~~-\left(g_{\rm eff}^{-}+g_{\rm eff}^{+}+i\hat{g}_{\rm eff}^{-}+i\hat{g}_{\rm eff}^{+}+2i\chi\right)C_{11},$ (43d)
$\dot{C}_{02}=-i\sqrt{2}\zeta^{*}C_{11}-2\left(g_{\rm eff}^{+}+i\hat{g}_{\rm eff}^{+}+i\chi\right)C_{02},$ (43e)
@@
Figure 4: Second order correlation function $\ln\left[g_{\rm out}^{(2)}(0)\right]$ influenced by Kerr nonlinearity when $g_{0}=1$GHz, $\kappa_{a}=0.2$GHz and $\zeta=0.5$GHz.
@@
Figure 5: The photon states in the resonator with $g_{0}=1$GHz, $\chi=4$GHz, $\kappa_{a}=0.2$GHz and $\zeta=0.5$GHz. As for the non-Markoivan parameter settings, $\omega_{a}L/c=\pi/3$ in (a) and (c), while $\omega_{a}L/c=\pi$ in (b) and (d). As for the classical driving field, $\epsilon=0.5$GHz in (a) and (b), while $\epsilon=4$GHz in (c) and (d).
To further clarify the photon number states in the microring resonator, the amplitudes of quantum states are simulated in Fig. 5. When the drive field is weak as $\epsilon=1$GHz, as in Fig. 5(a) with $\omega_{a}L/c=\pi/3$, the effective coupling between the waveguide and resonator is nonzero, which means that the photons in the resonator can be transmitted into the waveguide. However, because the driving field is weak and $\kappa_{a}>0$, there are almost no two-photon states in the resonator. When $\omega_{a}L=\pi$ and $\epsilon=1$GHz in Fig. 5(b), the effective coupling between the waveguide and resonator is zero according to Eq. (38), then the drive applied upon the resonator can compensate its decaying to the environment, then there can be two-photon states in the resonator. Based on this, when the drive field is stronger, as in Figs. 5(c,d) with $\epsilon=4$GHz, the populations for two-photon states can be larger, and photon blockade can occur due to the area in Fig. 4(b) with $\ln\left[g_{\rm out}^{(2)}(0)\right]<0$. For example, according to the steady numerical results in Fig. 5(c), $g_{\rm out}^{(2)}(0)\approx 0.4948<1$, and in Fig. 5(d), $g_{\rm out}^{(2)}(0)\approx 0.4961<1$. Besides, the comparison between Figs. 5(c,d) also illustrates that the non-Markovian interaction between resonator and waveguide evaluated by $\omega_{a}L$ can influence the steady photonic states in the microring resonator. When $\omega_{a}L/c=\pi$ in Fig. 5(d), the counterclockwise photon mode cannot be easily decayed to the waveguide, then the population for the steady two-photon states can be larger than that in Fig. 5(c).
As for more details on the experimental realizations of the proposal in this paper, according to the on-chip fabrications in Refs. [10, 30, 6, 11], the resonant frequency of the microring resonator can be 10 GHz to 100 GHz, the length of the waveguide constructing the coherent feedback loop can be 0.1 mm to 10 mm [10], then the non-Markovian proposal realized by microring resonators coupled to the waveguide at two distinct points can be fabricated on-chip with the length of several millimeters [10, 62, 70, 18, 49, 4, 73].
Based on the above parameter settings of microring resonators, the proposal in this paper can be further generalized to complex networks with multiple resonators. For example, a one-dimensional array of microring resonators can be coupled to a waveguide [67, 26, 10, 47], and the non-Markovian properties in such networks not only arise from the two-point coupling between one microring resonator and the waveguide as in this paper, but also arises from the transmission delays of photons between two microring resonators. Additionally, two-dimensional networks in the format of topological microring lattices [24, 63, 21, 38] can be used to construct coupled resonator optical waveguides. On one hand, in the one-dimensional and two-dimensional microring resonator networks, the edge states can be topologically-protected such that they are immune to defect scattering [1]. On the other hand, by coupling the topological microring resonator array to a waveguide, the topological property of the topological system (i.e., band structure, nonreciprocal property, and so on) can be detected and evaluated via the output end of the waveguide [67, 35, 27]. The generalization to one- and two-dimensional resonator networks can enhance the density of integration and provide new possible platforms for the realization of topological photonics with band structures and protect edge or entangled states [14, 25].